# Two-sample comparison through additive tree models for density ratios

### v2

@@ Abstract
The ratio of two densities characterizes their differences. We consider the two-sample comparison problem through estimating the density ratio given i.i.d. observations from each distribution. We propose additive tree models for the density ratio along with efficient algorithms for training these models using a new loss function called the balancing loss. With this loss, additive tree models for the density ratio can be trained using several algorithms originally designed for supervised learning. First, they can be trained from an optimization perspective through boosting algorithms including forward-stagewise fitting and gradient boosting. Moreover, due to the balancing loss function’s resemblance to an exponential family kernel, it can serve as a pseudo-likelihood for which conjugate priors exist, thereby enabling effective generalized Bayesian inference on the density ratio using backfitting samplers designed for Bayesian additive regression trees (BART). This generalized Bayesian strategy enables uncertainty quantification on the inferred density ratio, which is critical for applications involving high-dimensional and complex distributions, as in such problems uncertainty given limited data can often be substantial. We provide insights on the balancing loss through its close connection to the exponential loss in binary classification and to the variational form of $f$-divergence, in particular that of the squared Hellinger distance. Our numerical experiments demonstrate the accuracy and computational efficiency of the proposed approach while providing unique capabilities in uncertainty quantification. We demonstrate the application of our method in a case study involving assessing the quality of generative models for microbiome compositional data.
@@ 1 Introduction
For example, in large-scale biomedical studies involving the comparison of genetic measurements and various biomarkers between patients and controls, there is often no question that some types of two-sample differences do exist, and with a sufficiently large sample size, the null of no difference at all will be rejected. It is more helpful to specifically pinpoint what such differences are in order to help understand their biological meaning. Another timely application is the assessment of generative models. The key to assessing the quality of different models and algorithms for generating synthetic samples that imitate the real sampling distribution of the training data lies in deciphering the distributional difference between the real sampling distribution and that induced by the generative models, which is essentially a two-sample comparison problem. Some generative models, such as generative adversarial networks (GANs) (Goodfellow et al., 2014), essentially iteratively apply two-sample comparison and use the estimated misfit between the two distributions to sequentially improve the generative model. Two-sample comparison is also a quintessential problem in causal inference, in which consistent estimation of treatment effects often relies on correctly adjusting for the difference in the covariate distributions between the treated and the control groups.
@@
Our focus here is on carrying out two-sample comparison in a nonparametric fashion through estimation rather than hypothesis testing, using the density ratio function as the functional “statistic” of interest for characterizing two-sample differences. The density ratio fully summarizes how two distributions differ across the sample space, and thus allows one not only to carry out hypothesis testing but also to identify the nature of the difference as well. If, for example, the estimated density ratio deviates “significantly” away from 1 on parts of the sample space, it allows us to draw statistical conclusions regarding whether, where, and how the two distributions differ. At the same time, the notion of “significant” deviation will require some form of uncertainty quantification for the inferred density ratio, which will become available under our formulation.
@@
One may reasonably raise the concern that nonparametric density ratio learning is simply “too hard” for complex, high-dimensional data because density estimation itself in such settings is already a challenging problem. Why then would one attempt an even harder problem of learning the ratio of two densities? Though it might appear counter-intuitive, we argue that density ratio estimation (DRE) is in fact a much easier problem than density estimation in most modern applications. A density by definition is always a density ratio, with the denominator being the corresponding base measure against which the density is computed. The typical base measure, for example the Lebesgue measure or the uniform density in continuous cases, is generally chosen out of mathematical convenience and lack of prior knowledge, without taking any account of the dataset at hand (other than that the observations are continuous). This base measure generally does not resemble any sampling distribution in high-dimensional or otherwise complex problems. When good reliable knowledge of the sampling distribution does exist, one would be in the much simpler classical setting of a “goodness-of-fit” problem. In contrast, in most applications of two-sample comparison, the two distributions are often quite similar to each other and can serve as a “baseline of reference” for each other. As such, the density ratio between these two samples then is often much simpler than the underlying densities (i.e., their ratios with respect to the Lebesgue measure).
@@
We list several desiderata for learning the density ratio and discuss how we aim to achieve them. First, the model or function class representing the density ratio should be flexible enough to effectively approximate complex forms of the underlying truth. It should scale well with both sample size and in dimensionality. To these ends, we propose a class of density ratio learners based on additive tree ensembles (see e.g., Friedman et al., 2000; Hastie et al., 2009). Additive tree ensembles are a natural candidate for the density ratio problem, because of their success in classification problems. In fact, in the machine learning literature, a widely used technique to learn the density ratio between two samples is the so-called “density-ratio trick” (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022), which involves first fitting a binary classifier to the two samples, and then, based on the estimated posterior odds of the sample label, computing the density ratio. Additive tree boosting such as AdaBoost is among the most effective methods for binary classification.
In contrast to this popular perspective, we hold that DRE should not and does not need to involve the indirect route of inverting a binary classifier based on additive tree ensembles, but can instead be directly learned from the latter. As we will show, the resulting estimate are often substantially more accurate than what the density -ratio trick produces, especially when the sample size is unbalanced and/or the two-sample difference is local. In these scenarios, losses designed for binary classification are poor proxies for learning the density ratio.
@@
The second desideratum is to attain uncertainty quantification for the estimated density ratio, thereby allowing statistical inference to proceed. This aspect has mostly been ignored in the existing literature on DRE. One exception is Bayesian nonparametric models for the density ratio using a single tree model (Ma and Wong, 2011; Soriano and Ma, 2017). While they provide a means to examine posterior uncertainty in a fully Bayesian manner, piecewise-constant functions represented by single trees of controlled complexity are very restrictive and therefore inadequate for approximating complex density ratios. Additive trees substantially enrich the class of functions, but they do not allow a simple likelihood formulation for modeling density ratios, and consequently, fully Bayesian uncertainty quantification for this function class is challenging. To overcome this, we propose a generalized Bayesian approach (Bissiri et al., 2016) using the balancing loss to construct a pseudo-likelihood for this model class. Interestingly, we show that, due to the resemblance of the balancing loss to an exponential family kernel, there exists a natural conjugate prior for the model parameters under the pseudo-likelihood, which allows standard sampling strategies for BART models (Chipman et al., 2010) to be directly adopted for sampling the Gibbs posterior on the density ratio.
The remainder of the paper is organized as follows. In Section 2 we start by introducing the balancing loss for learning the density ratio and highlight its connection to the variational form of the squared Hellinger distance and to the exponential loss for binary classification. We then introduce an additive tree model for the (log) density ratio along with two boosting algorithms for fitting this model based on forward stage-wise fitting and gradient boosting, respectively. In Section 3, we introduce a generalized Bayesian inference framework that uses the balancing loss as a (log) pseudo-likelihood and show that one can accomplish Bayesian backfitting sampling on the model using standard algorithms for BART by exploiting the analogy of the pseudo-likelihood to an exponential family. In Section 4, we evaluate the performance of the proposed methods in comparison with the existing approaches in a few simulation settings, followed by a case study that assesses and comparing multiple generative models for microbiome compositions in Section 5. We give concluding remarks in Section 6.
2 Method
@@ 2.1 The balancing loss for density ratio learning
Let $r$ denote an estimate of the density ratio. In the following discussion, we assume that $r$ is positive $\mu$-almost everywhere. To introduce an algorithm for optimizing $r$, we define a loss function called the balancing loss based on the following fact: If $r$ is equal to the true density ratio $p/q$, then for any measurable subset $B\subset\Omega$, we have
$\int_{B}r^{-1/2}pd\mu=\int_{B}r^{1/2}qd\mu=\int_{B}\sqrt{pq}d\mu.$ (1)
When $B=\Omega$, the whole sample space, this corresponds to the affinity between $p$ and $q$. The first equation implies that if we weight the two samples respectively with $r^{-1/2}$ and $r^{1/2}$, their probability over any measurable set $B$ becomes equal. In this sense, we say that the two samples are “balanced”. Motivated by this observation, we introduce the notation $w=r^{1/2}$ to denote the balancing function, which is intended to approximate the square root of the density ratio $\sqrt{p/q}$. We then consider the following balancing loss function:
@@ 2.2 Connections to existing concepts in statistical learning
2.2.1 Connection 1: the “density ratio trick” based on binary classification
The first perspective is related to a common technique widely adopted in the machine learning community for estimating density ratios, often referred to as the “density-ratio trick”, which involves turning the problem into a supervised, binary classification task on the two group labels (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022). It involves inverting a binary classifier to compute the density ratio. We show that the balancing loss arises naturally in this context as a variant of the exponential loss widely used in this context, but only the balancing loss targets the density ratio directly and thus is more robust than the density ratio trick against scenarios such as when the two sample sizes are unbalanced and/or the two-sample differences exist locally in a small subset of the sample space. These will also be validated experimentally later in our numerical examples where a direct comparison with the density ratio trick will be provided.
The balancing loss can be seen as a modification of the exponential loss, which is used for classification with the AdaBoost, described in Friedman et al. (2000). To see this, we suppose that, in generating the two samples, one first randomly decides which sample to draw the observation from with equal probability. That is, one first draws a random variable $y$ that takes 1 and -1 with a 50-50 chance. Given $y$ one generates the corresponding sample $x$ from $P$ if $y=1$ and from $Q$ if y=1y=-1, respectively. Then it is straightforward to show that
$\mathbb{E}_{p}[w^{-1}]+\mathbb{E}_{q}[w]\propto\mathbb{E}[e^{-yF}],$
@@ 2.2.1 Connection 1: the “density ratio trick” based on binary classification
We note that the exponential loss used in AdaBoost in fact integrates $y$ with respect to its marginal distribution from the data rather than using a 50-50 split. Because the population minimizer for the AdaBoost is the log posterior odds between the two samples (Friedman et al., 2000), which is the product of the density ratio and the prior odds, one may wonder then why the original AdaBoost cannot suffice for learning density ratios since one can simply multiply by a constant (i.e., the ratio of the proportion of the two samples) to get an estimate of the density ratio. The main reason is that, by weighting the loss using the marginal proportion of the two samples, vanilla AdaBoost prioritizes classification accuracy over density ratio estimation; this can be most clearly demonstrated when there are parts of the sample space in which one of the samples is much more abundant than the other, which can occur when the sample sizes are very different or when the two sample densities have different supports or tail behaviors. Such cases are in fact quite common in multivariate applications, and numerical evaluations will be provided in Section 4.
@@
For a given convex function $f$, the general $f$-divergence is defined as $D_{f}(P\|Q):=\int f(\frac{p}{q})dQ$. It is known (Rényi, 1961; Csiszár, 1964; Ali and Silvey, 1966; Nguyen et al., 2010; Polyanskiy and Wu, 2025) that the dual form can be written as an optimization problem with respect to a functional class $\mathcal{F}$ mapping from $\mathcal{X}$ to $\mathbb{R}$,
$D_{f}(P\|Q)\geq\sup_{g\in\mathcal{F}}\int[gdP-f^{c}(g)dQ]$ (2)
where $f^{c}$ is the conjugate dual function of $f$, defined as $f^{c}:=\sup_{u\in\mathbb{R}}\{uv-f(u)\}$. The dual objective is referred to as the variational form of the $f$-divergence. In the case of the squared Hellinger divergence, $f(u)=\frac{1}{2}(\sqrt{u}-1)^{2}$, and the dual objective is $\sup_{g<\frac{1}{2}}\Big{\{}\mathbb{E}_{P}[g(X)]\;-\;\mathbb{E}_{Q}\!\Big{[}\frac{g(X)}{1-2g(X)}\Big{]}\Big{\}}$ with the optimum value $g^{*}=\frac{1}{2}-\frac{1}{2}\sqrt{\frac{q}{p}}$, based on a direct application of Lemma 1 in Nguyen et al. (2010). Now let $w=(1-2g)^{-1}$, the variational form for the squared Hellinger divergence becomes
$\sup_{w>0}\left\{1-\frac{1}{2}\int w^{-1}pd\mu-\frac{1}{2}\int wqd\mu\right\}$ (3)

### v3

@@ Abstract
The ratio of two densities provides a direct characterization of their differences. We consider the two-sample comparison problem by estimating this ratio given i.i.d. observations from two distributions. To this end, we propose additive tree models for density ratio estimation along with efficient algorithms using a new loss function, the balancing loss. The loss allows tree-based models to be trained using several algorithms originally designed for supervised learning, such as forward-stagewise optimization and gradient boosting. Moreover, the balancing loss resembles an exponential family kernel, and it can serve as a pseudo-likelihood with conjugate priors. This property enables generalized Bayesian inference on the density ratio using backfitting samplers designed for Bayesian additive regression trees (BART). Our Bayesian strategy provides uncertainty quantification for the inferred density ratio, which is critical for applications involving high-dimensional and data-limited distributions with potentially substantial uncertainty. We further show connections of the balancing loss to the exponential loss in binary classification and to the variational form of $f$-divergence, particularly the squared Hellinger distance. Numerical experiments demonstrate that our method achieves both accuracy and computational efficiency, while uniquely providing uncertainty quantification. Finally, we demonstrate its application to assessing the quality of generative models for microbiome compositional data.
Keywords: Generalized Bayesian inference, density ratio estimation, tree boosting, generative models
@@ 1 Introduction
For example, in large-scale biomedical studies involving the comparison of genetic measurements and various biomarkers between patients and controls, there is often no question that some types of two-sample differences do exist, and with a sufficiently large sample size, the null of no difference at all will be rejected. It is more helpful to specifically pinpoint what such differences are in order to help understand their biological implication. Another timely application is the assessment of generative models. The key to assessing the quality of different models and algorithms for generating synthetic samples that imitate the real sampling distribution of the training data lies in deciphering the distributional difference between the real sampling distribution and that induced by the generative models, which is essentially a two-sample comparison problem. Some generative models, such as generative adversarial networks (GANs) (Goodfellow et al., 2014), essentially iteratively apply two-sample comparison and use the estimated misfit between the two distributions to sequentially improve the generative model. Two-sample comparison is also a quintessential problem in causal inference, in which consistent estimation of treatment effects often relies on correctly adjusting for the difference in the covariate distributions between the treated and the control groups.
@@
Our focus here is on carrying out a two-sample comparison in a nonparametric fashion through estimation rather than hypothesis testing, using the density ratio function as the functional “statistic” of interest for characterizing two-sample differences. The density ratio fully summarizes how two distributions differ across the sample space, and thus allows one not only to carry out hypothesis testing but also to identify the nature of the difference as well. If, for example, the estimated density ratio deviates “significantly” away from 1 on parts of the sample space, it allows us to draw statistical conclusions regarding whether, where, and how the two distributions differ. At the same time, the notion of “significant” deviation will require some form of uncertainty quantification for the inferred density ratio, which will become available under our formulation.
@@
One may reasonably raise the concern that nonparametric density ratio learning is simply “too hard” for complex, high-dimensional data because density estimation itself in such settings is already a challenging problem. Why then would one attempt an even harder problem of learning the ratio of two densities? Though it might appear counterintuitive, we argue that density ratio estimation (DRE) is in fact a much easier problem than density estimation in most modern applications. A density by definition is always a density ratio, with the denominator being the corresponding base measure against which the density is computed. The typical base measure, for example, the Lebesgue measure or the uniform density in continuous cases, is generally chosen out of mathematical convenience and lack of prior knowledge, without taking any account of the dataset at hand (other than that the observations are continuous). This base measure generally does not resemble any sampling distribution in high-dimensional or otherwise complex problems. When good reliable knowledge of the sampling distribution does exist, one would be in the much simpler classical setting of a “goodness-of-fit” problem. In contrast, in most applications of two-sample comparison, the two distributions are often quite similar to each other and can serve as a “baseline of reference” for each other. As such, the density ratio between these two samples then is often much simpler than the underlying densities (i.e., their ratios with respect to the Lebesgue measure).
@@
We list several desiderata for learning the density ratio and discuss how we aim to achieve them. First, the model or function class representing the density ratio should be flexible enough to effectively approximate complex forms of the underlying truth. At the same time, it should scale well with both sample size and dimensionality. To these ends, we propose a class of density ratio learners based on additive tree ensembles (see e.g., Friedman et al., 2000; Hastie et al., 2009). Additive tree ensembles are a natural candidate for the density ratio problem, because of their success in classification problems. In fact, in the machine learning literature, a widely used technique to learn the density ratio between two samples is the so-called “density-ratio trick” (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022), which involves first fitting a binary classifier to the two samples, and then, based on the estimated posterior odds of the sample label, computing the density ratio. Additive tree boosting such as AdaBoost (Friedman et al., 2000) is among the most effective methods for binary classification.
In contrast to this popular perspective, we hold that DRE should not and does not need to involve the indirect route of inverting a binary classifier based on additive tree ensembles, but can instead be directly learned from the latter. To establish the direct route, we propose a new loss function called “balancing loss”, which has a functional form similar to the exponential loss for the classification AdaBoost (Friedman et al., 2000) but is minimized by the true density function. As we will show, the resulting estimates are often substantially more accurate than what the density-ratio trick produces, especially when the sample size is unbalanced, that is, when the setting is more challenging for classifiers that aim to maximize the total successful classification rate. We demonstrate the limitation of adopting the classification AdaBoost in the experiments.
@@
The second desideratum is to attain uncertainty quantification for the estimated density ratio, thereby allowing statistical inference to proceed. This aspect has mostly been ignored in the existing literature on DRE. One exception is Bayesian nonparametric models for the density ratio using a single tree model (Ma and Wong, 2011; Soriano and Ma, 2017). While they provide a means to examine posterior uncertainty in a fully Bayesian manner, piecewise-constant functions represented by single trees of controlled complexity are very restrictive and therefore inadequate for approximating complex density ratios. Also, more importantly, they introduce generative models for the two samples, and this approach can be inefficient in terms of the two-sample comparison, since estimating the common distributional structure is redundant. As such, generalizing these existing Bayesian models to build an additive tree model with which we can focus on the two-sample difference is quite challenging. To overcome this, we propose a generalized Bayesian approach (Bissiri et al., 2016) using the balancing loss to construct a pseudo-likelihood for this model class. Interestingly, we show that, due to the resemblance of the balancing loss to an exponential family kernel, there exists a natural conjugate prior for the model parameters under the pseudo-likelihood, which allows standard sampling strategies for BART models (Chipman et al., 2010) to be directly adopted for sampling the Gibbs posterior on the density ratio.
The remainder of the paper is organized as follows. In Section 2 we start by introducing the balancing loss for learning the density ratio and highlight its connection to the variational form of the squared Hellinger distance and to the exponential loss for binary classification. We then introduce an additive tree model for the (log) density ratio along with two boosting algorithms for fitting this model based on forward stage-wise fitting and gradient boosting, respectively. In Section 3, we introduce a generalized Bayesian inference framework that uses the balancing loss as a (log) pseudo-likelihood and show that one can accomplish Bayesian backfitting sampling on the model using standard algorithms for BART by exploiting the analogy of the pseudo-likelihood to an exponential family. In Section 4, we evaluate the performance of the proposed methods in comparison with the existing approaches in a few simulation settings, followed by a case study that assesses and compares multiple generative models for microbiome compositions in Section 5. We give concluding remarks in Section 6.
2 Boosting Algorithms for Density Ratio Estimation
In this section, we propose new boosting algorithms for density ratio estimation with a novel loss function called balancing loss. We first formulate the balancing loss and discuss its connection to other learning problems. Then, we introduce optimization procedures minimizing the loss, using the boosting algorithms (Friedman, 2001).
@@ 2.1 The Balancing Loss for Density Ratio Learning
Let $r$ denote an estimate of the density ratio. In the following discussion, we assume that $r$ is positive $\mu$-almost everywhere. To introduce an algorithm for optimizing $r$, we define a loss function called the balancing loss based on the following fact: If $r^{*}$ is equal to the true density ratio $p/q$, then for any measurable subset $B\subset\Omega$, we have
$\int_{B}(r^{*})^{-1/2}p\,d\mu=\int_{B}(r^{*})^{1/2}q\,d\mu=\int_{B}\sqrt{pq}\,d\mu.$ (1)
When $B=\Omega$, the whole sample space, this corresponds to the affinity between $p$ and $q$. The first equation implies that if we weight the two samples respectively with $(r^{*})^{-1/2}$ and $(r^{*})^{1/2}$, their probability over any measurable set $B$ becomes equal. In this sense, we say that the two samples are “balanced”. Motivated by this observation, we introduce the notation $w=r^{1/2}$ to denote the balancing function, which is intended to approximate the square root of the density ratio, $\sqrt{p/q}$. We then consider the following balancing loss function:
@@
Note that by the arithmetic-geometric mean inequality, the loss is bounded below by $2\int\sqrt{pq}\,d\mu$, and the equation holds if $w^{-1}p=wq$, or equivalently, $w=\sqrt{p/q}$. This simple justification is an advantage of selecting this functional form compared to other choices that can follow Eq. (1), such as $(\mathrm{E}_{p}[w^{-1}]-\mathrm{E}_{q}[w])^{2}$.
@@ 2.2 Connections to Existing Concepts in Statistical Learning
Connection 1: The “Density-ratio Trick” Based on Binary Classification
The first perspective is related to a common technique widely adopted in the machine learning community for estimating density ratios, often referred to as the “density-ratio trick”, which involves turning the problem into a supervised, binary classification task on the two group labels (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022). It involves inverting a binary classifier to compute the density ratio. As we discuss below, the balancing loss arises naturally as a variant of the exponential loss, which is used for classification with AdaBoost (Friedman et al., 2000), but only the balancing loss targets the density ratio directly and thus is more robust than the density ratio trick against scenarios such as when the two sample sizes are unbalanced and/or the two-sample differences exist locally in a small subset of the sample space. These will also be validated experimentally later in our numerical examples where a direct comparison with the density ratio trick will be provided.
To see the connection between the exponential loss and the proposed balancing loss, we suppose that, in generating the two samples, one first randomly decides which sample to draw the observation from with equal probability. That is, one first draws a random variable $y$ that takes 1 and -1 with a 50-50 chance. Given $y$ one generates the corresponding sample $x$ from $P$ if $y=1$ and from $Q$ if y=1y=-1, respectively. Then it is straightforward to show that
$\mathbb{E}_{p}[w^{-1}]+\mathbb{E}_{q}[w]\propto\mathbb{E}[\exp(-yF)],$
@@
We note that the exponential loss used in AdaBoost in fact integrates $y$ with respect to its marginal distribution from the data rather than using a 50-50 split. Because the population minimizer for AdaBoost is the log posterior odds between the two samples (Friedman et al., 2000), theoretically we can estimate the density ratio via the density ratio trick (Sugiyama et al., 2012), namely,
$\frac{p(x)}{q(x)}=\frac{\mathrm{Pr}(y=-1)}{\mathrm{Pr}(y=1)}\frac{\mathrm{Pr}(y=-1\mid x)}{\mathrm{Pr}(y=1\mid x)},$
where the prior ratio is replaced with the sample size ratio in practice. This approach, however, can introduce difficulties if the sample size is unbalanced. Under such imbalance, AdaBoost tends to show poor performance for classifying the observations with a small sample size, because the influence of the rare group under the empirical exponential loss $\sum_{i}e^{-y_{i}F(x_{i})}$ ($x_{i}$ and $y_{i}$ are an observation and its class, respectively) is substantially weak, which can result in a severe bias (see e.g., Sun et al., 2007, for a more in-depth discussion on this issue). Also, as a problem specific to density ratio estimation, the classification algorithm does not necessarily incorporate the information of prior odds explicitly, and multiplying the prior odds, that is, subtracting the information of the prior, can introduce a bias. An additional experiment for confirming the existence of the latter problem is provided in Supplementary Materials D.
The issues with the density-ratio trick caused by different sample sizes could be addressed by applying different weights to the two groups so that the two groups have equal weights in total. Similar ideas of re-weighting the minor/major groups are proposed in the context of classification for unbalanced data, as, for example, the cost-sensitive boosting (Sun et al., 2007). To the best of our knowledge, however, neither its application to density ratio estimation nor, more importantly, its extension for Bayesian inference is discussed in the literature. Moreover, we need to mention that an alternative calibration method of the estimated classification probability specialized for density ratio estimation is proposed in Cranmer et al. (2015), but the performance of AdaBoost is not substantially improved as we see in Section 4.
@@
For a given convex function $f:[0,\infty)\to(-\infty,\infty]$ with $f(1)=0$, and $P$ is absolutely continuous with respect to $Q$, the general $f$-divergence is defined as $D_{f}(P\|Q):=\int f(\frac{p}{q})\,dQ$. It is known (Rényi, 1961; Csiszár, 1964; Ali and Silvey, 1966; Nguyen et al., 2010; Polyanskiy and Wu, 2025) that the dual form can be written as an optimization problem with respect to a functional class $\mathcal{F}$ mapping from $\mathcal{X}$ to $\mathbb{R}$,
$D_{f}(P\|Q)\geq\sup_{g\in\mathcal{F}}\int[g\,dP-f_{c}(g)\,dQ],$
where $f_{c}$ is the conjugate dual function associated with $f$, defined as $f_{c}(v):=\sup_{u\in\mathbb{R}}\{uv-f(u)\}$. The equality holds when the subdifferential $\partial f(q_{0}/p_{0})$ contains an element in $\mathcal{F}$. The subdifferential of $f$ is defined as $\partial f(t):=\left\{z\in\mathbb{R}:f(s)\geq f(t)+z(s-t),\forall s\in\mathbb{R}\right\}$. See details in Lemma 1 of Nguyen et al. (2010). The dual objective is referred to as the variational form of the $f$-divergence. In the case of the squared Hellinger divergence, $f(u)=\frac{1}{2}(\sqrt{u}-1)^{2}$, and the dual objective is $\sup_{g<\frac{1}{2}}\Big\{\mathbb{E}_{P}[g(X)]\;-\;\mathbb{E}_{Q}\!\Big[\frac{g(X)}{1-2g(X)}\Big]\Big\}$ with the optimum value $g^{*}=\frac{1}{2}-\frac{1}{2}\sqrt{\frac{q}{p}}$, based on a direct application of Lemma 1 in Nguyen et al. (2010). Now let $w=(1-2g)^{-1}$, the variational form for the squared Hellinger divergence becomes
$\sup_{w>0}\left\{1-\frac{1}{2}\int w^{-1}p\,d\mu-\frac{1}{2}\int wq\,d\mu\right\}$